OpenStudy (anonymous):

Given the equation P2 = A3, what is the orbital period, in days, for the planet Venus? (Venus is located 0.72 AU from the sun?) 72 days 225 days 500 days 5,375 days

OpenStudy (fifciol):

using Keppler's 3rd law: \[\frac{ T_v^2 }{R_v^3 }=\frac{ T_E^2 }{R_E^3}=\frac{ T_v^2 }{ (0,72AU)^3 }=\frac{ (365,25)^2 }{ AU^2 }\rightarrow T_v= 225 d\]